The sum of three consecutive odd integers is -3. Find the value of the least of the three.
Found 2 solutions by greenestamps, mccravyedwin:
Answer by greenestamps(13327)



The average of three consecutive odd integers is the middle integer.

In this problem the sum of the three integers is -3, so the average is -3/3 = -1.

So the consecutive odd integers are -3, -1, and 1.

ANSWER: -3


Answer by mccravyedwin(421)

Let x = the value of the least of the three consecutive integers, then
x+1 = the value of the middle of the three consecutive integers. And
x+2 = the value of the largest of the three consecutive integers.

x + x+1 + x+2 = -3.  Solve for x.  That'll be the answer.

Edwin